If 100 students took a test, the class average was 76%. The girls averaged 70% and the boys averaged 80% how many girls participated in the test
Answer by TimothyLamb(4379)

---
x = boys
y = girls
---
x + y = 100
80x + 70y = 76*100
---
put the system of linear equations into standard form:
---
x + y = 100
80x + 70y = 7600
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = boys = 60
y = girls = 40
---
system matrix
1 1
80 70
---
inverse of system matrix
-7 0.1
8 -0.1
---
determinant of system matrix = -10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php